# Distance measurements from the internal dynamics of globular clusters: Application to the Sombrero galaxy (M 104)

### v1

@@ 5.1 Distance measurement with the GCVD relation
5.1.1 Distances from different samples
@@
We also matched -band photometry from the ground-based study of Rhode & Zepf (2004) with our sample, yielding 68 GCs. Rhode & Zepf (2004) adopted a single value for the extinction correction of , while Spitler et al. (2006) adopted . We therefore adjusted the Rhode & Zepf (2004) -band magnitudes by -0.008 mag to match the Spitler et al. (2006) and Harris et al. (2010) photometry.
@@ 5.1.2 Systematics
The effect of aperture and fibre centering on the velocity dispersion measurement is discussed in Sect. 3.2. These corrections translate to a uncertainty on our distances of 0.03 Mpc (0.2%) (i.e., this is the difference in distance we obtain with GCVD between the measured velocity dispersions and the aperture-corrected dispersions). Similarly, the impact of different photometry and photometric samples is discussed in Sect. 5.1.1, from which we obtain a standard deviation on the three different distances of 0.16 Mpc (1.7% distance uncertainty).
To investigate the impact of varying on GCVD, we used the sizes from Harris et al. (2010) where available. To also include the remaining sample, we assumed for the other GCs a mean size of 2.44 pc, the mean of the size measurements (see Sect. 4). We then performed bootstrap resampling with replacement, where we only selected GCs with to match the Milky Way and M 31 distributions. In all cases the GCVD slope was fixed to the fiducial value, and we just consider the impact on the offset . As a result of this exercise, we find that the offset can shift by up to 0.2 Mpc (2.2%).
@@
5.1.3 Dependence on GC parameters
After fitting the GCVD relation to the M 104 GCs, we find a scatter of the residuals of 6%. This is very similar to the scatter in the GC samples of the Milky Way, M 31, and Centaurus A (see Fig. 5) and suggests that this scatter is intrinsic to such GC samples. In Fig. 6, we colour-coded the GCs according to different parameters to test if the scatter correlates with GC parameters.
In the left panel of Fig. 6, the GCs are colour-coded according to their metallicity. The metallicity values are drawn from the full M 104 sample described in Fahrion et al. (2025). Where available, we used metallicities from optical VLT/MUSE or GTC/OSIRIS spectroscopy, and for the remaining GCs the metallicities were derived directly from full spectrum fitting to the FLAMES spectra. While the GCs with measured dispersions cover a broad range of metallicities, there seems to be no clear correlation with the scatter in the magnitude-dispersion plane. Splitting the MW and M 31 GCs into metal-poor and metal-rich GCs, Beasley et al. (2024) found no difference in the GCVD relation’s slope or offset, in agreement with our finding for M 104.
The fourth panel shows GCs with available size measurements. There might be a hint that smaller-sized GCs lie to the right of the GCVD relation, but this is driven by three objects and a larger sample is required to test this dependence. However, the observed offset of SUCD1 could be explained by the size difference as it is a particularly large object with a reported half-light radius of pc (Hau et al., 2009), significantly larger than typical GCs (mean pc, Harris et al. 2010; Masters et al. 2010) As the GCVD relation is a consequence of the Virial Theorem, based on the relation between total mass and the combination of velocity dispersion squared and size (), the size can influence the position of a source on the relation, with the effect expected to scale linearly. With the much larger size of SUCD1, this effect might play a role, while the narrow range of GC sizes does not have a large impact.
The final panel in Fig. 6 shows the GCs colour-coded by their dynamical mass-to-light ratio as described in the following section. GCs with larger mass-to-light ratios appear to lie preferentially to the right of the GCVD relation. Given that the dynamical mass scales with the square of the velocity dispersion, this is perhaps not surprising, but indicates that the GCVD distance method might only by applicable for objects within a narrow range of mass-to-light ratios. Interestingly, SUCD1 is an outlier in this respect as it lies above the relation, but still has a larger mass-to-light ratio than most of the GCs (see Sect. 4). This is likely a consequence of the elevated mass-to-light ratio of SUCD1 stemming from its large size, while the higher mass-to-light ratios of the GCs are driven by an increased velocity dispersion.
@@ 6 Comparison to literature distances
As a well studied galaxy, multiple distance measurements to M 104 exist with varying accuracy. As described in Sect. 1 and discussed by McQuinn et al. (2016), distances to M 104 vary between 7 and 22 Mpc, with the more accurate measurements placing the galaxy between 8 and 10 Mpc (e.g. Jensen et al. 2003; Spitler et al. 2006; McQuinn et al. 2016). McQuinn et al. (2016) note that distance measurements for M 104 based on the Tully-Fisher relation generally find larger distances, between 11 – 22 Mpc (Bottinelli et al., 1984; Tully & Fisher, 1988; Sorce et al., 2014). This may be due to the peculiar morphology and spatially restricted gas distribution of M104 that make the morphological classification of this galaxy somewhat uncertain, and set it apart from more typical elliptical and spiral galaxies.
In Fig. 7, we show distances from the literature in comparison with our fiducial GCVD distance of Mpc from fitting the combined -band magnitudes (see Sect. 5). As can be seen from this figure, the GCVD distance compares well against literature measurements from various sources. For example, the median of the shown GCLF, PNLF, SBF, and TRGB distances is at 9.25 Mpc, and calculating the weighted mean taking the statistical uncertainties into account, we find Mpc, with a dispersion of 0.60 Mpc.
@@
While most distances agree with our GCVD distance within 1 of the combined statistical uncertainties, there are exceptions. For example, the GCLF distance of Mpc reported by Spitler et al. (2006) would place the Sombrero galaxy significantly closer than our GCVD distance. As Spitler et al. (2006) discuss, assuming a distance of 9.0 Mpc would make the turnover magnitude of the GCLF about 0.3 mag brighter than the universally adopted value. However, investigation of early-type galaxies in the Virgo and Fornax galaxy clusters have found that massive galaxies often have brighter GCLF turnover magnitudes (by 0.1 - 0.3 mag compared to low-mass galaxies, Villegas et al. 2010). The bright turnover magnitude in M 104 might therefore be a consequence of M 104 being a massive galaxy ( Jardel et al. 2011; Karunakaran et al. 2020).
@@ 7 Conclusions
Using -band magnitudes from space- and ground-based observations of 77 GCs, we applied the GCVD relation with fixed slope and obtained a distance of =9.00 0.29 (stat.) 0.26 (sys.) Mpc. We find an intrinsic scatter of 6%. Testing different photometric samples including -band magnitudes derived from other filters finds consistent distance measurements.
@@ Acknowledgments
KF acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101103830. ACS acknowledges support from FAPERGS (grants 23/2551-0001832-2 and 24/2551-0001548-5), CNPq (grants 314301/2021-6, 312940/2025-4, 445231/2024-6, and 404233/2024-4), and CAPES (grant 88887.004427/2024-00). A.F.M acknowledges support from RYC2021-031099-I of MICIN/AEI/10.13039/501100011033/ UE NextGenerationEU/PRTR. O.M. is grateful to the Swiss National Science Foundation for financial support under the grant number PZ00P2_202104. This work made use of Astropy:333http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2013, 2018, 2022). Based on observations collected at the European Southern Observatory under ESO programme 114.274W (PI: Fahrion).
Appendix A Table of dispersion measurements

### v2

@@ 5.1 Distance measurement with the GCVD relation
We excluded SUCD1 from all distance measurements due to its peculiar properties. Additionally, it is still debated if UCDs build the high-mass end of the genuine GC population (e.g. Bekki et al. 2002; Mieske et al. 2002), or are rather the remnant nuclear star clusters (NSCs) of disrupted dwarf galaxies (e.g. Bekki et al. 2003; Pfeffer and Baumgardt 2013). While they likely are a mixed population (e.g. Norris et al. 2014; Wang et al. 2023), with its large size and high mass, SUCD1 is likely a NSC-type UCD (Hau et al., 2009). The stripping process (Pfeffer et al., 2016) as well as the complex stellar populations found in many NSCs (e.g. Kacharov et al. 2018; Fahrion et al. 2021, 2024) might explain its offset from the GCVD relation.
5.2 Distances from different samples
@@
We also matched -band photometry from the ground-based study of Rhode and Zepf (2004) with our sample, yielding 68 GCs. Rhode and Zepf (2004) adopted a single value for the extinction correction of , while Spitler et al. (2006) adopted . We therefore adjusted the Rhode and Zepf (2004) -band magnitudes by -0.008 mag to match the Spitler et al. (2006) and Harris et al. (2010) photometry.
@@ 5.3 Systematics
The effect of aperture and fibre centring on the velocity dispersion measurement is discussed in Sect. 3.2. These corrections translate to an uncertainty on our distances of 0.03 Mpc (0.2%) (i.e., this is the difference in distance we obtain with GCVD between the measured velocity dispersions and the aperture-corrected dispersions). Similarly, the impact of different photometry and photometric samples is discussed in Sect. 5.2, from which we obtain a standard deviation on the three different distances of 0.16 Mpc (1.7% distance uncertainty).
To investigate the impact of varying on GCVD, we used the sizes from Harris et al. (2010) where available. To also include the remaining sample, we assumed for the other GCs a mean size of 2.44 pc, the mean of the size measurements. We then performed bootstrap resampling with replacement, where we only selected GCs with to match the Milky Way and M31 distributions. In all cases the GCVD slope was fixed to the fiducial value, and we just consider the impact on the offset . As a result of this exercise, we find that the offset can shift by up to 0.2 Mpc (2.2%).
@@
We also explored the validity of our assumption of using a fixed slope for the GCVD relation obtained from the Milky Way and M31 data. We ran the MCMC fitting on the Sombrero data leaving both the zeropoint and the slope in Eq. 3 free. As in Beasley et al. (2024) we chose uniform priors in the range (zeropoint), and (slope). For the full sample, we found . This would translate to a distance, D = Mpc. This slope is steeper than that of the fiducial Milky Way + M31 sample () at the level. If we only select Sombrero GCs in the same magnitude range as the fiducial relation (i.e., mag) we find and obtain D = Mpc which is consistent with our distance measurement with fixed slope. To conclude, there is some evidence that the slope for the brighter Sombrero GCs might be steeper than in the Milky Way and M31 sample, but for the bulk of the sample the assumption of a fixed slope based on fiducial relation is likely reasonable and larger samples across broad magnitude ranges are required to test the effects of the slope properly.
5.4 Dependence on GC parameters
After fitting the GCVD relation to the M104 GCs, we find a scatter of the residuals of 6%. This is very similar to the scatter in the GC samples of the Milky Way, M31, and Centaurus A (see Fig. 7) and suggests that this scatter is intrinsic to such GC samples. In Fig. 8, we show the orthogonal residuals from the GCVD fit as a function of different GC parameters (similar to Fig. 4. The orthogonal residuals refer to distance in and to be best-fit line, normalised to the best-fit line.
As can be seen in this figure, the residuals show no dependence on most of the GC parameters such as metallicity, colour, line-of-sight velocity, projected radius or mass-to-light ratio. However, as also noted in Fig. 4, there is a weak relation with the size and consequently the relaxation time.
@@ 6 Comparison to literature distances
As a well studied galaxy, multiple distance measurements to M104 exist with varying accuracy. As described in Sect. 1 and discussed by McQuinn et al. (2016), distances to M104 vary between 7 and 22 Mpc, with the more accurate measurements placing the galaxy between 8 and 10 Mpc (e.g. Jensen et al. 2003; Spitler et al. 2006; McQuinn et al. 2016). McQuinn et al. (2016) note that distance measurements for M104 based on the Tully-Fisher relation generally find larger distances, between 11 – 22 Mpc (Bottinelli et al., 1984; Tully and Fisher, 1988; Sorce et al., 2014). This may be due to the peculiar morphology and spatially restricted gas distribution of M104 that make the morphological classification of this galaxy somewhat uncertain, and set it apart from more typical elliptical and spiral galaxies.
In Fig. 9, we show distances from the literature in comparison with our fiducial GCVD distance of (stat.) 0.26 (syst.) Mpc from fitting the combined -band magnitudes (see Sect. 5). We show both the statistical uncertainty and the quadrature sum of statistical and systematic uncertainty (in light grey) to visualise a total uncertainty. Distances based on the Tully-Fisher relation were excluded from Fig. 9 due to the aforementioned reasons. As can be seen from this figure, the GCVD distance compares well against literature measurements from various sources. For example, the median of the shown GCLF, PNLF, SBF, and TRGB distances is at 9.25 Mpc, and calculating the weighted mean taking the statistical uncertainties into account, we find Mpc, with a dispersion of 0.60 Mpc.
@@
While most distances agree with our GCVD distance within 1 of the combined statistical uncertainties, there are exceptions. For example, the GCLF distance of Mpc reported by Spitler et al. (2006) would place the Sombrero galaxy significantly closer than our GCVD distance. As Spitler et al. (2006) discuss, assuming a distance of 9.0 Mpc would make the turnover magnitude of the GCLF about 0.3 mag brighter than the universally adopted value. However, investigation of early-type galaxies in the Virgo and Fornax galaxy clusters have found that massive galaxies often have brighter GCLF turnover magnitudes (by 0.1 - 0.3 mag compared to low-mass galaxies, Villegas et al. 2010). The bright turnover magnitude in M104 might therefore be a consequence of M104 being a massive galaxy (; Jardel et al. 2011; Karunakaran et al. 2020).
@@ 7 Conclusions
Using -band magnitudes from space- and ground-based observations of 77 GCs, we applied the GCVD relation with fixed slope and obtained a distance of (stat.) 0.26 (sys.) Mpc. We find an intrinsic scatter of 6%. Testing different photometric samples including -band magnitudes derived from other filters finds consistent distance measurements.
@@ Acknowledgments
We thank the two referees for constructive reports that have helped to improve this manuscript. KF thanks Francisco Aros for helpful discussions. KF acknowledges funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101103830. ACS acknowledges support from FAPERGS (grants 23/2551-0001832-2 and 24/2551-0001548-5), CNPq (grants 314301/2021-6, 312940/2025-4, 445231/2024-6, and 404233/2024-4), and CAPES (grant 88887.004427/2024-00). A.F.M acknowledges support from RYC2021-031099-I of MICIN/AEI/10.13039/501100011033/ UE NextGenerationEU/PRTR. O.M. is grateful to the Swiss National Science Foundation for financial support under the grant number PZ00P2_202104. KF, MB, AFM, ACS, GvV, AG acknowledge support from PID2024-16088NB-I00. This work made use of Astropy:333http://www.astropy.org a community-developed core Python package and an ecosystem of tools and resources for astronomy (Astropy Collaboration et al., 2013, 2018, 2022). Based on observations collected at the European Southern Observatory under ESO programme 114.274W (PI: Fahrion).
Appendix A Radial trends of magnitudes and dispersions
Figure 10 shows the -band magnitudes and velocity dispersions of the M104 GCs as a function of projected distance from the centre, illustrating that some of the brightest GCs in the sample are also among the brightest. The close match in the versus trend stems from the GCVD relation and we do not see any evidence for inflated velocity dispersions due to contamination from the galaxy light in the fiber apertures.
Appendix B Table of dispersion measurements